# Privacy in Speech Technology

### v2

@@ V-B Subjective evaluation of privacy
In speaker recognition, computers are now on par or better than humans  [hansen2015speaker] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [vestman2020methods, vestman2020voice]. We can thus expect that computers are better at evaluating privacy in well-defined tasks. However, there are at least two caveats. First, in identifying private information, human listeners can potentially use properties or features of speech signals other than those computers use. Subjective listeners can thus potentially discover new weaknesses that computers miss. Second, in scenarios where humans are the potential attackers extracting information (as in threats 3, 5, and 7 in fig. 1), the true measure of privacy is quantified by subjective listening. Thus, it is often prudent to complement objective privacy metrics with subjective tests.
@@
In the subjective evaluation of privacy, the listening task should be defined to match the threat. For example, if the threat is re-identification, subjects should be tasked with identifying subjects from anonymized speech signals. Similarly, if the threat is speakers’ attribute inference, the subjects should evaluate the extent to which they can identify such attributes. For a practical implementation of such subjective evaluation of privacy, see [tomashenko2022voiceprivacy].
V-C Objective performance in downstream tasks
@@
Metrics used for quantifying utility depend on the application. In applications where a computer analyzes the speech signal, we quantify utility by measuring the accuracy of analysis. For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [rohit2024asr]. By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy-protections impede utility.
V-D Subjective performance in downstream tasks
@@
If a human user is to listen to the output signal, then that user is the ultimate measure of performance [zacharov2018sensory]. For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [vincent2018audio]. Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [P800, MUSHRA]. Similarly, if privacy protections degrade a speech signal, we can readily quantify such degradations by listening tests.
@@
Subjective listening tests are, however, expensive, time-consuming, and noisy. Individual subjects, especially those without training (known as naïve subjects), are typically unable to repeat a sequence of tests such that results would match. Subjects are also slow since they need to listen to samples one at a time, at real-time speeds, require time to answer between each sample, a multitude of subjects need to do the same test to gain statistical confidence in results, and tests need oversight by trained personnel. Automated tests, known as objective measures of perceptual quality, are an alternative to such subjective tests. They estimate the outcome of subjective listening tests by computational means. For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [rix2001perceptual]. Later improvements include POLQA and ViSQOL [beerends2013perceptual, chinen2020visqol]. Such objective measures can be used to automate quality testing at a lower price and faster speed so that repeated estimates with the same input give the same result. Observe that researchers should carefully review the specifications of such algorithms to determine whether they apply to the problem at hand to ensure reliable results.
V-E Privacy-preserving systems
Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy. Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level. Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speakers’ identity. New samples are compared to such biometric references to find matching attributes. The three key concepts in such evaluation are [ISOIEC224745_2022, nautsch2019preserving]
  • Renewability – multiple, unlikable biometric references can be created from the same biometric samples and used to identify the person without revealing information about the biometric samples.
Together, unlinkability and irreversibility thus minimize the amount of information leaked if biometric references are revealed. At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them.
VI Perception and Psychology of Privacy
@@
People tend to have a deep sense of ownership of some things, both material and immaterial. In particular, people have a feeling of ownership toward information about themselves [cichy2014extending, dawkins2017psychological]. Such feelings are, to some extent, detached from the material consequences of breaches in privacy. For example, even if publishing an audio recording revealing a secret intimate relationship would not have direct economic consequences, it can cause psychological damage and harm personal relationships. The effect of threats to privacy then necessarily becomes a question of psychology and social sciences. To qualitatively or quantitatively measure such effects, we also need subjective user tests.
@@
Concurrently, people have well-established social rules regarding human-to-human privacy [petronio2002boundaries]. Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [nass1993anthropomorphism, cornelius2021acceptance], such that they will likely assume that devices apply human-like social rules to privacy. Observe that such social rules are related to human-like behavior and performance. That does not reveal whether they are applicable to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past. How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [thompson2021machine]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
VI-A Perception and Experience of Privacy
@@
Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [pepino2020distrust]. Clearly, users do not like using services they perceive as threatening. Understanding how people perceive privacy is interesting in its own right, but such understanding is essential in the design of effective user interfaces.
@@
Perception and experience of privacy can be approached in two alternative ways. We can make user studies where people interact with either machines or each other [leschanowsky2024review]. The distinction is important in the sense that human-to-human interaction relies on social rules that are well-established and developed over a long time. We can thus expect them to be stable over time, whereas human-computer interaction is continuously evolving as people learn more. Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance but can probably not rely on observing effects related to super-human performance. Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [leschanowsky2024review].
@@
Studies of human-computer interaction thus characterize the desired phenomenon directly. The compromise is, however, that people’s understanding of privacy with devices might not reflect the true level of privacy, and conclusions made based on the users’ opinions might then not reflect their true preferences and those preferences are moving targets. As people gain more experience with and as they learn more about technology, their attitudes, and perceptions of it change [leschanowsky2024review].
@@
Despite these shortcomings, both types of experiments are essential for improving our understanding of privacy and for improving technology. For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [zarazaga19_interspeech, das2021thesis, zarazaga2022thesis, backstrom2021intuitive, Leschanowsky2020] and multiple-bed patient rooms in hospitals have privacy-concerns [reinten2017speech]. Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [bruggemeier2022perceptions], voice interfaces with unknown features cause fear in users, and reduces retention of services [yeasmin2020privacy, Lau2018, lutz2021privacy] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [Kurz2021].
VI-B User Interface Design
@@
For usable privacy, it is important that information about privacy is readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [bruggemeier2022perceptions, so2021voice, lewis1998designing, Lau2018, Kurz2021, dhiya2023privacy, iachello2007end, sivaraman2018smart]. As a parallel to speech interfaces, visual interfaces can use lights or icons for monitoring and tactile interfaces for controlling the level of privacy. While sound can be used to monitor system status with sonification [hildebrandt2016continuous, iber2021auditory], this is not in widespread use within signaling of privacy. It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [pulkki2015communication]. To be effective, the sound should be perceivable when the user consciously wants to check the privacy level and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape that users consciously register such changes. Furthermore, user interfaces should give correct information about and allow controlling the privacy level [Lau2018, backstrom2020privacy].
@@
In any case, it is imperative that services are designed such that they reflect the true extent of privacy. Observe that it is likely possible to design systems that communicate an advanced level of privacy even when the system does not respect user privacy. In fact, service providers may have short-term incentives to follow such approaches to design as long as they improve overall user satisfaction. However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [willis2020deception]. Through deception, such dark patterns lull users into believing they are safe when, in fact, they are abused. Therefore, good design of privacy should actively communicate and enable control of the true extent of privacy. Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [Kurz2021].
VI-C Privacy and Trust
@@
The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [leschanowsky2024review]. In particular, if users experience that a service is detrimental to their privacy or, more generally, to their well-being, they will obviously not trust it. We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [balsa2022cryptography].
@@
While the meaning of trust may seem intuitively clear, it is however difficult to define exactly. A classic definition is [deutch1973resolution]
@@
This definition highlights the psychological nature of trust [simpson2007psychological]. From the current engineering perspective, this translates to the statistical confidence that the outcome of a process is useful. Here, usefulness refers to what the user, subjectively perceives as useful. Implicitly, this furthermore requires that we predict the outcome of that process in the form of a statistical distribution of possible outcomes and that the expectation over the possible outcomes is that the process is useful. Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [Chen2003, Xie2009].
@@
It is useful and ethical to design speech services that are trustworthy. This means that systems should be such that it would be logical for users to trust them [hardin1996trustworthiness]. Speech technology should thus be designed to have useful outcomes with statistically significant confidence, especially with respect to privacy.
A practical approach for reducing the likelihood of negative outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [ometov2018multi]. In fig. 4, this can be implemented by applying layers of privacy protection in each successive connection and device as well as in their internal modules and components, as described in section IV-E. Activation of an undesired functionality of the system or leaking private information would thus require that every subsequent module fails simultaneously, which is unlikely in most scenarios.
@@
VII Content Categories and Applications
Recognizing the speaker’s identity then becomes the natural starting point for studies in privacy. We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g. [chouchane21_interspeech, granqvist20_interspeech, mawalim20_interspeech, mawalim20_interspeech, nautsch19b_interspeech, nelus19_interspeech, nelus19b_interspeech, noe21_interspeech, perero2022x, zewoudie2021federated, zewoudie2021use, teixeira22_interspeech, treiber2019privacy, nelus19_interspeech, nelus19b_interspeech, BAI202165, han2020voice, KINNUNEN201012, hansen2015speaker, parthasarathi2013wordless, pathak11_interspeech, parthasarathi11_interspeech, pathak2012ppsa, jimenez2017two, portelo2014bprivacy]. Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g. [noe20_interspeech, pierre22_interspeech, prajapati21_interspeech, shamsabadi2023diff, srivastava20_interspeech, Srivastava2022taslp, mawalim20_interspeech, maouche20_interspeech, Fang2019]. Speaker characterization is the natural complement of speaker idenfication [nautsch2019preserving, nguyen2010automatic]. Such methods related to speaker identity can be applied for example to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming. Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research. With respect to privacy, it contains two obvious challenges for privacy. We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g. [shoemate2022sottovoce, nguyen2023federated, aloufi21_interspeech, garg20b_interspeech, shah21_interspeech, srivastava19_interspeech, tomashenko2022privacy], or we can use natural language processing to anonymize the text content, e.g. [liu22_interspeech, maouche2022enhancing, adelani20_interspeech, novotney21_interspeech, mahendran2021privacy, habernal2021differential].
Privacy is even more important in always-on applications like wake-word detection, i.e. when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”. This application is more sensitive exactly because it is always on as users cannot choose when data is processed [koppelmann21_interspeech, koppelmann22_interspeech, stoidis22_interspeech, yang21_interspeech]. The always-on characteristic is also prominent in assisted-living applications, which can ”monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [haider19_interspeech]. Such ambient voice interfaces are often implemented through acoustic sensor networks which pose their own challenges [koppelmann22_interspeech, oconnor20_interspeech, malekzadeh2018replacement, zewoudie2021use, zarazaga2020acoustic, nelus2021].
Speech enhancement refers to removing background noises and distortions from the desired speech signal (private message) [benesty2008springer]. This task can have an impact on privacy in two ways. First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [zarazaga21_interspeech, rech2022multi]. Attenuating background noises and competing speakers, as well as removing reverberation, can thus improve privacy. Again, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [nelus2021, zarazaga2022thesis, hendriks2013ppdistributed].
Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [adler2006towards, nautsch20_interspeech, nautsch2019preserving, ISObiometric, edpb2021guidelines]. Some, however, limit the term biometric to refer only to physical and behavioral characteristics. Such properties include health [natarajan2021prioris], emotions [feng22b_interspeech] and gender identity [nelus2018itg, stoidis21_interspeech, stoidis22_interspeech, nelus19_interspeech], each warranting their own treatment.
@@ VIII Legal and Societal Landscape
While we have not (yet) seen breaches related to speech technology with consequences of comparable magnitude, these parallels highlight the potential effect of breaches. Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [Lynskey2019GuardianPrivacyScandal, hern2019, day2019].
@@
Section VI-B also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy. We can easily find examples where users do not have real choices in protecting their privacy. For example, suppose all friends and family of a user use a particular platform for social and voice interaction even if its privacy configuration is inadequate. The individual user faces then the choice of either disconnecting from their social network or compromising privacy. This applies to all users with a sufficiently large portion of their communities participating in that platform. This is a variation of the prisoner’s dilemma, where no single user has sufficient incentive to transition to a solution that would be best for everyone.
@@
On the individual level, the examples of potential exploits in table II cause clear damage to individuals, including psychological harm in particular (e.g. [natarajan2021prioris]). Additional dangers include stalkers [lu2019abusers, hulsen2022stalkers, avast2021stalkerware]. While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [levy2020intimate].
@@
These examples demonstrate the inherent need for society-level regulation of speech technology with respect to privacy. Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [eu2016gdpr, nautsch19c_interspeech], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [california2018ccpa]. While these laws cover only a small percentage of the global population, as cloud services typically operate globally, they need the capability to follow local laws. In many cases, it can be easiest to apply the strictest laws on all users, such that the most strict laws benefit the privacy of all users. Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a large impact also on users outside the scope of these regulations.
@@
With respect to regulation, an important consequence of the objective measures of privacy in section V is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence. This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [eu2016gdpr], which does not explicitly leave room for statistical uncertainty. This is reflected for example in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [edpb2021guidelines, page 13, §31]
@@
This leaves the interpretation open. It is possible to argue that it is never possible to obtain absolute confidence in speaker identification such that the GDPR is never triggered. It is also possible to argue that all voice data contains personal information that can be used to uniquely identify a person, such that all processing must have a valid legal basis. Both interpretations lead to absurdity, which suggests that the truth must lie somewhere in the middle. In fact, the GDPR in practice requires (see [edpb2021guidelines, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy. Authors of the GDPR are thus clearly aware that it is impossible to give absolute guarantees of privacy but that the impact assessment (i.e., objective measures of privacy) must necessarily be based on statistical measures, even if such measures have not been defined.
However, even in the absence of legal requirements, the current author would encourage all researchers, developers, data analysts, project managers, and policymakers to perform impact assessments in data projects, data management, and data policies, anytime when there is or could be the slightest concern that the dataset imposes risks to privacy. A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [franzke2021data].
While governments are in the process of regulating privacy, corporations and non-governmental organizations have also realized that proper privacy is an opportunity. For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [ovn2023] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [mydata2023, poikola2020mydata]. Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [spsc2023]. It is as far as we know, the world-wide largest community focused on this topic.

### v3

@@ VI-B Subjective evaluation of privacy
In speaker recognition, computers are now on par or better than humans [36] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [176, 177]. We can thus expect that computers are better at evaluating privacy in well-defined tasks. However, there are at least two caveats. First, in identifying private information, human listeners can potentially use properties or features of speech signals other than those computers use. Subjective listeners can thus potentially discover new weaknesses that computers miss (though the opposite is also true). Second, in scenarios where humans are the potential attackers extracting information (as in threats 3, 5, and 7 in fig. 1), the true measure of privacy is quantified by subjective listening. Thus, it is often prudent to complement objective privacy metrics with subjective tests.
@@
In the subjective evaluation of privacy, the listening task should be defined to match the threat. For example, if the threat is re-identification, subjects should be tasked with identifying subjects from anonymized speech signals. Similarly, if the threat is speakers’ attribute inference, the subjects should evaluate the extent to which they can identify such attributes.
As an example, in the VoicePrivacy 2020 challenge, subjective speaker verifiability was measured by asking subjects to rate, on a scale of 1 to 10110, the similarity between an anonymized target and a reference speaker, where the reference was either the same or a different speaker [98]. If the distributions of similarity scores have significant differences, then the anonymization has failed. Alternatives to the 1 to 10110 Likert-scale include, for example, AB (are these samples the same person?) and ABX (which of AB is equal to X?) tests. For a comprehensive exposition of subjective evaluation methods, see [178].
VI-C Objective performance in downstream tasks
@@
Metrics used for quantifying utility depend on the application. In applications where a computer analyzes the speech signal, we quantify utility by measuring the accuracy of analysis. For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [179]. By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy-protections impede utility.
VI-D Subjective performance in downstream tasks
@@
If a human user is to listen to the output signal, then that user is the ultimate measure of performance [178]. For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [180]. Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [181, 182]. Similarly, if privacy protections degrade a speech signal, we can readily quantify such degradations by listening tests.
@@
Subjective listening tests are, however, expensive, time-consuming, and noisy. Individual subjects, especially those without training (known as naïve subjects), are typically unable to repeat a sequence of tests such that results would match. Subjects are also slow since they need to listen to samples one at a time, at real-time speeds, require time to answer between each sample, a multitude of subjects need to do the same test to gain statistical confidence in results, and tests need oversight by trained personnel. Automated tests, known as objective measures of perceptual quality, are an alternative to such subjective tests. They estimate the outcome of subjective listening tests by computational means. For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [183]. Later improvements include POLQA and ViSQOL [184, 185]. Such objective measures can be used to automate quality testing at a lower price and faster speed so that repeated estimates with the same input give the same result. Observe that researchers should carefully review the specifications of such algorithms to determine whether they apply to the problem at hand to ensure reliable results.
VII Perception and psychology of privacy
Much of the work of the speech community is focused on technical aspects of privacy-preserving algorithms. Yet speech devices are used by and designed for people, and they use people’s voice data. It is therefore essential to know what people think of speech devices, how they use and want to use them, and their preferences about privacy. Much of this work falls under the fields of psychology, user-interface design, and social sciences. This section is an introduction to these topics.
VII-A Psychological privacy and ownership
@@
People tend to have a deep sense of ownership of some things, both material and immaterial. In particular, people have a feeling of ownership toward information about themselves [186, 187]. Such feelings are, to some extent, detached from the material consequences of breaches in privacy. For example, even if publishing an audio recording revealing a secret intimate relationship would not have direct economic consequences, it can cause psychological damage and harm personal relationships. The effect of threats to privacy then necessarily becomes a question of psychology and social sciences. To qualitatively or quantitatively measure such effects, we also need subjective user tests.
@@
Concurrently, people have well-established social rules regarding human-to-human privacy [25]. Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [188, 189], such that they will likely assume that devices apply human-like social rules to privacy. Observe that such social rules are related to human-like behavior and performance. That does not reveal whether they are applicable to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past. How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [190]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
VII-B Perception and experience of privacy
@@
Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [191]. Clearly, users do not like using services they perceive as threatening. Understanding how people perceive privacy is interesting in its own right, but such understanding is essential in the design of effective user interfaces.
@@
Perception and experience of privacy can be approached in two alternative ways. We can make user studies where people interact with either machines or each other [192]. The distinction is important in the sense that human-to-human interaction relies on social rules that are well-established and developed over a long time. We can thus expect them to be stable over time, whereas human-computer interaction is continuously evolving as people learn more. Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance but can probably not rely on observing effects related to super-human performance. Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [192].
@@
Studies of human-computer interaction thus characterize the desired phenomenon directly. The compromise is, however, that people’s understanding of privacy with devices might not reflect the true level of privacy, and conclusions made based on the users’ opinions might then not reflect their true preferences and those preferences are moving targets. As people gain more experience with and as they learn more about technology, their attitudes, and perceptions of it change [192].
@@
Despite these shortcomings, both types of experiments are essential for improving our understanding of privacy and for improving technology. For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [193, 17, 18, 194, 195] and multiple-bed patient rooms in hospitals have privacy-concerns [196]. Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [197], voice interfaces with unknown features cause fear in users, and reduces retention of services [198, 199, 200] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [201].
VII-C User interface design
@@
For usable privacy, it is important that information about privacy is readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [197, 202, 203, 199, 201, 204, 205, 206]. As a parallel to speech interfaces, visual interfaces can use lights or icons for monitoring and tactile interfaces for controlling the level of privacy. While sound can be used to monitor system status with sonification [207, 208], this is not in widespread use within signaling of privacy. It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [209]. To be effective, the sound should be perceivable when the user consciously wants to check the privacy level and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape that users consciously register such changes. Furthermore, user interfaces should give correct information about and allow controlling the privacy level [199, 15].
@@
In any case, it is imperative that services are designed such that they reflect the true extent of privacy. Observe that it is likely possible to design systems that communicate an advanced level of privacy even when the system does not respect user privacy. In fact, service providers may have short-term incentives to follow such approaches to design as long as they improve overall user satisfaction. However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [210]. Through deception, such dark patterns lull users into believing they are safe when, in fact, they are abused. Therefore, good design of privacy should actively communicate and enable control of the true extent of privacy. Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [201].
VII-D Privacy and trust
@@
The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [192]. In particular, if users experience that a service is detrimental to their privacy or, more generally, to their well-being, they will obviously not trust it. We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [211].
@@
While the meaning of trust may seem intuitively clear, it is however difficult to define exactly. A classic definition is [212]
@@
This definition highlights the psychological nature of trust [213]. From the current engineering perspective, this translates to the statistical confidence that the outcome of a process is useful. Here, usefulness refers to what the user, subjectively perceives as useful. Implicitly, this furthermore requires that we predict the outcome of that process in the form of a statistical distribution of possible outcomes and that the expectation over the possible outcomes is that the process is useful. Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [214, 215].
@@
It is useful and ethical to design speech services that are trustworthy. This means that systems should be such that it would be logical for users to trust them [216]. Speech technology should thus be designed to have useful outcomes with statistically significant confidence, especially with respect to privacy.
A practical approach for reducing the likelihood of negative outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [217]. In fig. 4, this can be implemented by applying layers of privacy protection in each successive connection and device as well as in their internal modules and components, as described in section V-E. Activation of an undesired functionality of the system or leaking private information would thus require that every subsequent module fails simultaneously, which is unlikely in most scenarios.
@@ VIII Legal and societal landscape
While we have not (yet) seen breaches related to speech technology with consequences of comparable magnitude, these parallels highlight the potential effect of breaches. Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [5, 6, 221].
@@
Section VII-C also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy. We can easily find examples where users do not have real choices in protecting their privacy. For example, suppose all friends and family of a user use a particular platform for social and voice interaction even if its privacy configuration is inadequate [222]. The individual user faces then the choice of either disconnecting from their social network or compromising privacy. This applies to all users with a sufficiently large portion of their communities participating in that platform. This is a variation of the prisoner’s dilemma, where no single user has sufficient incentive to transition to a solution that would be best for everyone.
@@
On the individual level, the examples of potential exploits in table II cause clear damage to individuals, including psychological harm in particular (e.g. [90]). Additional dangers include stalkers [223, 224, 225] and oppression [222]. While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [8].
@@
These examples demonstrate the inherent need for society-level regulation of speech technology with respect to privacy. Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [226, 227], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [228]. While these laws cover only a small percentage of the global population, as cloud services typically operate globally, they need the capability to follow local laws. In many cases, it can be easiest to apply the strictest laws on all users, such that the most strict laws benefit the privacy of all users. Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a large impact also on users outside the scope of these regulations. More recently, other countries have followed suit, for example through the “Personal Information Protection Law of the People’s Republic of China” [229] and “Lei Geral de Proteção de Dados” (Brazilian General Data Protection Law, LGPD) [230].
@@
With respect to regulation, an important consequence of the objective measures of privacy in section VI is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence. This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [226], which does not explicitly leave room for statistical uncertainty. This is reflected for example in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [89, page 13, §31]
@@
This leaves the interpretation open. It is possible to argue that it is never possible to obtain absolute confidence in speaker identification such that the GDPR is never triggered. It is also possible to argue that all voice data contains personal information that can be used to uniquely identify a person, such that all processing must have a valid legal basis. Both interpretations lead to absurdity, which suggests that the truth must lie somewhere in the middle. In fact, the GDPR in practice requires (see [89, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy. Authors of the GDPR are thus clearly aware that it is impossible to give absolute guarantees of privacy but that the impact assessment (i.e., objective measures of privacy) must necessarily be based on statistical measures, even if such measures have not been defined. Fortunately, first mathematical characterizations of the legal criteria have started to appear [169, 170].
However, even in the absence of legal requirements, the current author would encourage all researchers, developers, data analysts, project managers, and policymakers to perform impact assessments in data projects, data management, and data policies, anytime when there is or could be the slightest concern that the dataset imposes risks to privacy. A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [231].
While governments are in the process of regulating privacy, corporations and non-governmental organizations have also realized that proper privacy is an opportunity. For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [232] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [233, 234]. Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [235]. It is as far as we know, the world-wide largest community focused on this topic.